Given the parent functions f(x) = 5x - 1 and g(x) = 3x - 9, what is g(x) - f(x)?
Answer by fcabanski(1391)

g(x) - f(x) = 3x - 9 -(5x - 1) = 3x - 9 -5x + 1 = -2x - 8